student 1 completed two more than 6 times as many homework as student 2. together they completed 51 homework. how many homework did each student complete? i try to solve it but i got a fraction answer. i don like this. 2+6x=51. pls show me where do i get wrong. thanks!
Answer by kvijioct(5)

Assume that Student 2 completed 'x' homework questions.
So, Student 1 completed (6x+2) homework questions.
Together, they solved 51 questions.
So, x+(6x+2)=51
7x+2=51
7x=51-2
7x=49
x=49/7
x=7
So, Student 2 completed 7 questions.
Student 1 completed (6x+2)= {(6*7)+2}
=42+2
=44 questions.